archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details We re going to update the firmware of the modems SDSL TG605 from version 8 2 3 10 to version 8 2 7 7 this Thursday September 12th from 01 00 to 06 00 During the update the modem will relaunch on the new version of the firmware causing a few minutes cut off from the internet Date Thursday

    Original URL path: http://status.ovh.es/?do=details&id=5374 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    ADSL SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details From 0 00 tonight we will be updating the ADSL devices of NRA BDV75 BEA75 PRO75 and VOL75 30 minutes downtime is expect following the intervention Date

    Original URL path: http://status.ovh.es/?do=details&id=5340 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    optic measurements Date Monday 02 September 2013 20 04PM Reason for closing Done Comments 3 Comment by OVH Monday 02 September 2013 14 26PM We have carried out the measures on our side We are no longer getting a signal on an optic strand between SPR and MPL A ticket has been opened with our fibre provider Comment by OVH Monday 02 September 2013 18 10PM The cut in the

    Original URL path: http://status.ovh.es/?do=details&id=5311 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    a second probleme concerning BGP SFR have not told us the correct networks on this port We re keeping the port down and we re working with SFR to resolve the issue The northern SFR collect traffic is currently flowing via Paris Date Friday 23 August 2013 00 14AM Reason for closing Done Comments 1 Comment by OVH Friday 23 August 2013 00 14AM SFR told us a bit about

    Original URL path: http://status.ovh.es/?do=details&id=5261 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details Tonight at around 1 00am we will be shutting down the national and North region FT collect port due to optic migration Service outage of less than 10mins is to be expected Date Wednesday 21 August 2013 02

    Original URL path: http://status.ovh.es/?do=details&id=5251 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Friday August 9th at 14 00 Date Tuesday 20 August 2013 15 51PM Reason for closing Done Comments 7 Comment by OVH Friday 09 August 2013 15 54PM We are starting the maintenance works we are inserting the 10G cards on the relevant sites Comment by OVH Friday 09 August 2013 15 55PM We are working on the first side of the loop Comment by OVH Friday 09 August 2013 18 47PM We have an optic attenuation issue We are halting the intervention Optimisation works are required on the optic loop We will start again on Monday 12th Securisation is operational in case of power failure Comment by OVH Monday 12 August 2013 16 23PM We are continuing with the intervention on the Lille optic loop We are going to reconfigure the EDFA amps in order to move the loop onto 10G Comment by OVH Monday 12 August 2013 17 12PM We are moving the amp to boi59 securisation is operational in case of power failure Comment by OVH Friday 16 August 2013 13 07PM We have succeeded to raise one side of the 10G loop however we are still having reception issues with securisation We are going to intervene

    Original URL path: http://status.ovh.es/?do=details&id=5198 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details Tonight around midnight we ll cut the collect FT port of Bordeaux in preparation for optical switch tomorrow Subscribers will be all switched to Roubaix Ins Date Tuesday 13 August 2013 12 09PM Reason for closing Done

    Original URL path: http://status.ovh.es/?do=details&id=5216 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details We detected a fiber cut between the NRA of HEL59 and Roubaix A ticket was opened investigations are in progress on their side There is no service interruption security was resumed Date Monday 12 August 2013 09 57AM

    Original URL path: http://status.ovh.es/?do=details&id=5207 (2015-10-15)
    Open archived version from archive



  •